DIRC Operations Report
Period from 2008/03/19 to 2008/04/09
Oncall/Fast-Mon: Jose Benitez, Jennifer Prendki .
See Run 7 integrated luminosity here
PEP, BaBar activities/issues since last DIRC meeting
Finished taking data on the Y(2S) and scanned the region above the Y(4S) up to 11.2GeV.
After that rescanned Y(6S) with more statistics.
Increased current in the LER upto 3200mA and other machine studies for super-B.
No major problems on the BaBar side.
DIRC activities/issues since last DIRC meeting
Current detector status (Cosmics run from last night):
- Smooth data taking until the end.
- Monday night after final beams were dumped at midnight, a global calibration was taken then cosmics wer taken all night. Yesterday morning a random trigger run was taken and a long-ADC calibration was taken.
- HV crates are now off.
some tails on the timing distributions (DFB noise?).
Lumi and Background History
- What to do about Wienner crates? Turn off --> Chiller should be turned off --> Should chiller water be drained?
- Same as before. see:DIRC Spares